# Junyu Cao and Mohsen Bayati

### v1

@@ 2 Model
In this section, we introduce our model setup. We first introduce the supervised learning oracle that will be potentially used during the initial learning stage, and then present the fine-tuning procedure through human comparisons during the refinement stage.
@@ 2.1 Supervised Learning and Underspecifications
Given data $\{(X_{i},Y_{i})\}_{i\leq n}$ where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$, the goal is to learn the true function that is parameterized by $\boldsymbol{\vartheta}^{*}\subseteq\mathbb{R}^{d}$, $h_{\boldsymbol{\vartheta}^{*}}\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{H}$ is the function class. Each data point is generated from a stochastic response:
@@
where $\epsilon_{i}$ is the noise with mean 0 and variance $\sigma^{2}$, where $\sigma$ could potentially be very large.
We specifically consider a high-dimensional problem that can be represented by low-dimensional manifolds; that is,
$h_{\boldsymbol{\vartheta}^{*}}(x)=f_{\boldsymbol{\theta}^{*},\varphi(\cdot)}(x)\,,$
where $f(\cdot)$ is assumed to be a generalized linear function in $\boldsymbol{\theta}^{*}$ and $\varphi(x)$. Here $\boldsymbol{\theta}^{*}\in\Theta\in\mathbb{R}^{s}$ is the unknown parameter [^2] We denote the parameter vector using the bold symbol., $\varphi(x)\in\mathbb{R}^{s}$ is the low-dimensional representation of feature $x$. It is assumed that $s\ll d$. For simplicity, we will remove the subscript $\varphi(\cdot)$ from $f$ when it is clear from the context.
We first provide the following two examples as an illustration.
@@
The parameter $\boldsymbol{\vartheta}^{*}$ has nonzero entries in the first $s$ dimensions and has zero entries in the rest of dimensions. In the linear model, $f_{\boldsymbol{\vartheta}^{*}}(x)=\sum_{i=1}^{s}\vartheta^{*}_{i}x_{i}$. In this case, $\boldsymbol{\theta}^{*}=(\vartheta^{*}_{1},\cdots,\vartheta^{*}_{s})$, $\varphi(x)=(x_{1},\cdots,x_{s})$, and $f(\cdot)$ is a linear function.
@@
In the generalized low-rank models, it is assumed that
$f_{\boldsymbol{\Theta}^{*}}(X)=\mu(\left\langle\boldsymbol{\Theta}^{*},xz^{\top}\right\rangle)=\mu(x^{\top}\boldsymbol{\Theta}^{*}z)$
where $\mu(\cdot)$ is the link function, $\boldsymbol{\Theta}^{*}\in\mathbb{R}^{d\times d}$ is a low-rank matrix and $X=xz^{\top}$, i.e., $\boldsymbol{\Theta}^{*}=\sum_{i=1}^{s}\sigma_{i}u_{i}v_{i}^{\top}$ where $u_{i}$ and $v_{i}\in\mathbb{R}^{d}$. Thus, $f_{\boldsymbol{\Theta}^{*}}(X)=\sum_{i=1}^{s}\sigma_{i}(x^{\top}u_{i})(v_{i}^{\top}z)$. In this case, $\varphi(X)$ is the projection of $X$ onto the low-rank space and $\boldsymbol{\theta}^{*}$ is the $s$-dimensional vector of singular values.
Suppose $h_{\boldsymbol{\vartheta}}$ can be estimated through a statistical loss function $\ell:\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}$, and the total loss is
@@
For example, $\ell$ can represent the squared loss, the regularized squared loss, among others. To deal with the high-dimensional issue, the algorithm learns the low-dimensional representation $\varphi(x)$ as well as the parameter $\boldsymbol{\theta}^{*}$. To learn the sparse linear models for example, one can apply lasso to select important features and learn the model parameters. To learn the low-rank models, one can use trace regression to learn the low-dimensional representations.
Define $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ as the function contained in $\mathcal{H}$ that minimizes empirical risk $L_{n}$. In the existing literature, it is very common that the decision-maker directly employs the predictor $h_{\hat{\boldsymbol{\vartheta}}_{n}}$. However, some issues may arise. In high-dimensional problems where significant noise is present, approaches striving to balance the variance-bias trade-off may occasionally diverge from intuitive expectations. We use the following example to illustrate this potential issue.
@@
Consider a linear pricing model with high-dimensional contexts, i.e., $Y_{i}=\vartheta_{1}\cdot p_{i}+\vartheta_{2}^{\top}X_{i}+\epsilon_{i}$ where $Y_{i}$ is the demand, $p_{i}$ is the price, and the vector $X_{i}$ represents features of product $i$, and also assume that $\vartheta_{1}<0$. However, in a high-noise regime we may incorrectly estimate a positive value for $\vartheta_{1}$. Please see Appendix 9.1 for numerical illustrations.
In Example 2.3, the positive correlation between price and demand is misinterpreted due to model underspecification. The requirement of $\vartheta_{1}<0$ should be specified based on human knowledge. In the face of the challenge posed by underspecification within a potentially very noisy environment, we take a critical step towards refinement, which we term “fine-tuning through human comparisons.” Human can do comparisons with a much higher precision. For example, based on the human knowledge, it can be immediately corrected that the demand and the price is negatively correlated. This particular stage comes into play after the output generated by our machine learning pipelines, represented as $\hat{\varphi}(\cdot)$ and $\hat{\boldsymbol{\theta}}_{0}$ (which denotes the estimator of $\varphi$ and $\boldsymbol{\theta}$), has been produced. During this stage, we focus on enhancing the model within a lower-dimensional space. We move on to the refinement step after the first stage where the kernel function, which provides a low-dimensional representation, has been effectively trained on the noisy-labeled data.
@@ 2.2 Utility Model of Human Comparisons
In this section, we introduce the human comparison model. Given two potential models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the system asks human to compare and select the better prediction model. If $f_{\boldsymbol{\theta}_{1}}$ is more preferable than $f_{\boldsymbol{\theta}_{2}}$, we denote it as $f_{\boldsymbol{\theta}_{1}}\succeq f_{\boldsymbol{\theta}_{2}}$.
The preference over the model $f_{\boldsymbol{\theta}}$ is captured by a utility function $u(\boldsymbol{\theta})$, where $u(\boldsymbol{\theta})$ depicts the distance between the model parameterized by $\boldsymbol{\theta}$ and the true model. That is, we assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$ where $d(\cdot,\cdot)$ is a distance function. The larger value of utility indicates a better model. The true model parameter $\boldsymbol{\theta}^{*}$ achieves the highest utility. For example, it can be the negative two-norm distance, i.e., $u(\boldsymbol{\theta})=-\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\|_{2}$. We formally state the assumption about the utility function as follows:
@@
[Utility function] Assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function. The utility function is unimodal in each coordinate. That is, for each coordinate $\theta_{i}$, when fixing all other coordinates, $u(\boldsymbol{\theta})$ is monotonically increasing when $\theta_{i}<\theta_{i}^{*}$ and monotonically decreasing when $\theta_{i}>\theta_{i}^{*}$. Moreover, the utility gets lower when $\boldsymbol{\theta}$ gets far away from the true parameter. The utility function can be either bounded or unbounded (e.g., the Euclidean distance). We give another example of the bounded utility function as an illustration.
@@
The comparison between $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ provides information on which side is closer to the true parameter $\boldsymbol{\theta}^{*}$. Define the hyperplane $H(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})=d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$, and half-space $H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})<d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$ and $H^{+}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})>d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$. The condition $u(\boldsymbol{\theta}_{1})\geq u(\boldsymbol{\theta}_{2})$ implies that $\boldsymbol{\theta}^{*}\in H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$. Assumption 2.2 implies that $H(\theta_{1},\theta_{2})$ contains a unique point when $\theta_{1}$ and $\theta_{2}$ are in a one-dimensional space.
@@
In practice, the comparison between two models can be conducted in various ways. For example, it can be evaluated on a single point with high confidence (either from human’s past experience or their observation), denoted as $\{x,y\}$. When comparing two models, the one with smaller square loss is more preferred, i.e.,
We will elaborate this process in Section 5.
The overall goal is to find $\boldsymbol{\theta}$ that achieves near the highest utility, or equivalently, gets close to the true model parameter $\boldsymbol{\theta}^{*}$ within a certain precision. Suppose that human’s selection process follows a random utility model (RUM). In the process of querying, human’s opinion regarding $f_{\boldsymbol{\theta}}$ is
$U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\beta_{h}+\xi,$
where $\beta_{h}$ represents the human bias, and $\xi$ is the unobserved noise. The human bias $\beta_{h}$ can vary between individuals. We assume the noise $\xi$ follows Gumbel distribution with parameter $\gamma$, which represents the expert level of the human. We say the human makes the correct choice if he selects $\boldsymbol{\theta}_{1}$ ($\boldsymbol{\theta}_{2}$) when $u(\boldsymbol{\theta}_{1})$ ($u(\boldsymbol{\theta}_{2})$) is higher than $u(\boldsymbol{\theta}_{2})$ ($u(\boldsymbol{\theta}_{1})$). A smaller value of $\gamma$ corresponds to a higher certainty of the selection correctness. When $\gamma$ approaches 0, human can always select the better one with almost 100% accuracy.
@@
For any two models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the probability that human will make the right selection is
@@
which is strictly larger than 1/2.
Lemma 2.6 implies that under RUM, the selection accuracy is always higher than 50%, even with the nonzero human bias. Intuitively, the differentiation would become harder when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ both get closer to the true parameter. However, how fast it converges to 1/2 when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ approach $\boldsymbol{\theta}^{*}$ depend on the distance function. Under this probabilistic selection model, we will design the algorithm and investigate the sample complexity. The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then extend to the multi-dimensional scenario.

### v2

@@ 3 Model
In this section, we introduce our model setup. We first introduce the supervised learning oracle that potentially is used during the initial learning stage, and then we present the fine-tuning procedure involving human feedback during the refinement stage. The two alternative ways of collecting human feedback are either to collect human labels directly or to ask humans to compare two answers. The former approach, which continues the label collecting procedure following the initial learning stage, is called pure supervised learning. As previously noted, we call the latter approach the Supervised Learning+Learning from Human Feedback framework.
@@ 3.1 Supervised Learning and Underspecifications
Given the data $\{(X_{i},Y_{i})\}_{i\leq n}$, where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$, the goal is to learn the true function that is parameterized by $\boldsymbol{\vartheta}^{*}\subseteq\mathbb{R}^{d}$, $h_{\boldsymbol{\vartheta}^{*}}\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{H}$ is the function class. Each data point is generated from a stochastic response:
@@
where $\epsilon_{i}$ is the noise with mean 0 and variance $\sigma^{2}$, and where $\sigma$ could potentially be very large.
We specifically consider a high-dimensional problem that can be represented by the low-dimensional generalized linear model; that is,
$h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle)\,,$
where $f(\cdot)$ is a link function. Here, $\boldsymbol{\theta}^{*}\in\Theta\in\mathbb{R}^{s}$ is the unknown parameter [^1] We denote the parameter vector using the bold symbol.; and $\varphi(x)\in\mathbb{R}^{s}$ is the kernel function, which is the low-dimensional representation of feature $x$. Our goal is to learn both the parameter $\boldsymbol{\theta}^{*}$ and representation $\varphi(x)$. We assume that $s\ll d$.
We provide the following two examples as illustrations.
@@
Suppose $x$, $\boldsymbol{\vartheta}^{*}\in\mathbb{R}^{d}$, and $h_{\boldsymbol{\vartheta}^{*}}(x)=\left\langle\boldsymbol{\vartheta}^{*},x\right\rangle$. The parameter $\boldsymbol{\vartheta}^{*}$ has nonzero entries in the first $s$ dimensions and has zero entries in the rest of the dimensions. In this case, $h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle)$ where $\boldsymbol{\theta}^{*}=(\vartheta^{*}_{1},\cdots,\vartheta^{*}_{s})$, $\varphi(x)=(x_{1},\cdots,x_{s})$, and $f(\cdot)$ is an identity function.
@@
Suppose $x\in\mathbb{R}^{d\times d}$. In the generalized low-rank models, we assume that $h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\Theta}^{*},x\right\rangle)$ where $\boldsymbol{\Theta}^{*}\in\mathbb{R}^{d\times d}$ is a low-rank matrix (i.e., $\boldsymbol{\Theta}^{*}=\sum_{i=1}^{s}\sigma_{i}u_{i}v_{i}^{\top}$ where $u_{i}$ and $v_{i}\in\mathbb{R}^{d})$; the trace inner product is defined as $\left\langle\boldsymbol{\Theta}^{*},x\right\rangle=tr(\boldsymbol{\Theta}^{*}x^{\top})$. Thus,
$h_{\boldsymbol{\vartheta}^{*}}(x)=f(\sum_{i=1}^{s}\sigma_{i}\cdot tr(u_{i}v_{i}^{\top}x^{\top}))=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle),$
where $\varphi(X)=[tr(u_{i}v_{i}^{\top}x^{\top})]_{i=1}^{s}$ is the projection of $x$ onto the low-rank space, and $\boldsymbol{\theta}^{*}=(\sigma_{1},\cdots,\sigma_{s})$ is the $s$-dimensional vector of singular values.
Suppose that $h_{\boldsymbol{\vartheta}}$ can be estimated through a statistical loss function $\ell:\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}$ and that the total loss is
@@
For example, $\ell$ can represent the squared loss and the regularized squared loss, among others. To deal with the high-dimensional issue, the algorithm learns the low-dimensional representation $\varphi(x)$, as well as the parameter $\boldsymbol{\theta}^{*}$. For example, to learn the sparse linear models, one can apply Lasso to select important features and to learn the model parameters. To learn the low-rank models, one can use trace regression to learn the low-dimensional representations.
Define $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ as the function contained in $\mathcal{H}$ that minimizes empirical risk $L_{n}$. In the existing literature, the decision maker very commonly employs the predictor $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ directly. However, some issues may arise. In high-dimensional problems, where significant noise is present, approaches striving to balance the variance–bias trade-off may violate fundamental principles, such as the natural expectation of negative demand elasticity in economics. We use the following example to illustrate this potential issue.
@@
Consider a linear pricing model with high-dimensional contexts – that is, $Y_{i}=\vartheta_{1}\cdot p_{i}+\vartheta_{2}^{\top}X_{i}+\epsilon_{i}$, where $Y_{i}$ is the demand, $p_{i}$ is the price, and the vector $X_{i}$ represents features of product $i$. Also, assume that $\vartheta_{1}<0$. However, in a high-noise regime, we may incorrectly estimate a positive value for $\vartheta_{1}$. Please see Appendix 10 for numerical illustrations.
In Example 3.3, the negative coefficient of the price in the demand function is misinterpreted because of model underspecification. The requirement of $\vartheta_{1}<0$ should be specified based on human knowledge. In the face of the challenge that underspecification poses in a potentially very noisy environment, we take a critical step toward refinement; we call this step “model alignment through human comparisons.” Humans can do comparisons with much greater precision. For example, on the basis of their knowledge, humans can immediately correct the negative correlation between demand and price. This particular stage comes into play after our machine learning pipelines generate and produce their output $\hat{\varphi}(\cdot)$ and $\hat{\boldsymbol{\theta}}_{0}$ (which denotes the estimator of $\varphi$ and $\boldsymbol{\theta}$) by training on the noisy-labeled data. During the second stage, we refine the model in a lower dimensional space. Throughout the paper, we focus our analysis on sparse linear models.
@@ 3.2 Utility Model of Human Comparisons
In this section, we introduce the human comparison model. Given two potential models, $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the system asks humans to compare and select the better prediction model; we call this step the model-level comparison. If $f_{\boldsymbol{\theta}_{1}}$ is preferable to $f_{\boldsymbol{\theta}_{2}}$, we denote the result as $f_{\boldsymbol{\theta}_{1}}\succeq f_{\boldsymbol{\theta}_{2}}$.
Remark 3.4 (Model-level Comparison and Sample-level Comparison)
In practice, the comparison between two models can be conducted in various ways. For example, the models can be evaluated on a single data point (either from the humans’ past experience or their observation), denoted as $\{x,y\}$. When comparing two models in this case, the one with a smaller squared loss is preferred; that is:
Alternatively, models also can be assessed on a batch of samples. We call this approach the sample-level comparison. Note that a model-level comparison is a broader concept. We elaborate on how to translate the model-level comparison to the sample-level comparison in Section 6. In this section, our focus is on a framework for undertaking a more general model-level comparison.
The preference over the model $f_{\boldsymbol{\theta}}$ is captured by a utility function $u(\boldsymbol{\theta})$, where $u(\boldsymbol{\theta})$ depicts the distance between the model parameterized by $\boldsymbol{\theta}$ and the true model. That is, we assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function. The larger utility value indicates the better model. The true model parameter $\boldsymbol{\theta}^{*}$ achieves the highest utility. For example, it can be the negative two-norm distance – that is, $u(\boldsymbol{\theta})=-\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\|_{2}$. We formally state the assumption about the utility function as follows:
@@
[Utility function] Assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function. The utility function is unimodal in each coordinate. That is, for each coordinate $\theta_{i}$, when fixing all other coordinates, $u(\boldsymbol{\theta})$ is monotonically increasing when $\theta_{i}<\theta_{i}^{*}$ and is monotonically decreasing when $\theta_{i}>\theta_{i}^{*}$. Moreover, the utility gets lower when $\boldsymbol{\theta}$ gets far away from the true parameter. The utility function can be either bounded or unbounded (e.g., the Euclidean distance). We give another example of the bounded utility function as an illustration.
@@
The comparison between $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ provides information on which side is closer to the true parameter $\boldsymbol{\theta}^{*}$. Define the hyperplane $H(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})=d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$, as well as half-space $H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})<d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$ and half-space $H^{+}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})>d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$. The condition $u(\boldsymbol{\theta}_{1})\geq u(\boldsymbol{\theta}_{2})$ implies that $\boldsymbol{\theta}^{*}\in H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$. Assumption 3.4 implies that $H(\theta_{1},\theta_{2})$ contains a unique point when $\theta_{1}$ and $\theta_{2}$ are in a one-dimensional space.
The overall goal is to find the $\boldsymbol{\theta}$ that achieves near the highest utility or that, equivalently, gets close to the true model parameter $\boldsymbol{\theta}^{*}$ within a certain precision. Suppose that the human selection process follows a random utility model (RUM). In the process of querying, the human’s opinion regarding $f_{\boldsymbol{\theta}}$ is
$U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\xi,$ (3.1)
where $\xi$ is the unobserved noise. We assume that the noise $\xi$ follows a Gumbel distribution with parameter $\gamma$, which represents the expert level of the human. We say human evaluators make the correct choice if they select $\boldsymbol{\theta}_{1}$ ($\boldsymbol{\theta}_{2}$) when $u(\boldsymbol{\theta}_{1})$ ($u(\boldsymbol{\theta}_{2})$) is higher than $u(\boldsymbol{\theta}_{2})$ ($u(\boldsymbol{\theta}_{1})$). A smaller value of $\gamma$ corresponds to a higher certainty of the selection’s correctness. When $\gamma$ approaches 0, humans can select the better one with almost 100% accuracy.
@@
Our model can also incorporate the human bias in the following way: $U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\beta_{h}+\xi$, where the human bias $\beta_{h}$ can vary between individuals. For simplicity, we focus on the utility model (3.1).
@@
For any two models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the probability that a human will make the right selection is
@@
which is strictly greater than 1/2.
Lemma 3.7 implies that under a RUM, the selection accuracy is always higher than 50%, even with the nonzero human bias. Intuitively, the differentiation would become harder when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ both get closer to the true parameter. However, how fast it converges to 1/2 when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ approach $\boldsymbol{\theta}^{*}$ depends on the distance function. Under this probabilistic selection model, we design the algorithm and investigate the sample complexity. The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then we extend to the multi-dimensional scenario.